OpenStudy (anonymous):

Help with PRE ALGEBRA problem: Find the x-intercept and the y-intercept for the graph of each equation. y = x - 1

OpenStudy (anonymous):

x intercept: when y is 0. 0=x-1 y intercept: when x is 0. y=0-1

OpenStudy (anonymous):

How did you get that? Sorry I forgot my notes at school. :(

OpenStudy (anonymous):

It is a rule

OpenStudy (anonymous):

I tried it and I got (1, 0 ) and ( 0 , -1 )

OpenStudy (anonymous):

Is that right @hedyeh99 ?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Yay! Can you help me with the next one? It's a bit harder

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

X = 9

OpenStudy (anonymous):

@hedyeh99

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

It just says x = 9 I got ( 0 , 9 ) for the x int and then none for y int

OpenStudy (anonymous):

It is going to be actually (9,0) for the x intercept

OpenStudy (anonymous):

Because when x is 9, y is 0

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Could you help with a few more? Sorry IM bad at this

OpenStudy (anonymous):

It's ok, I'll help you

OpenStudy (anonymous):

y + 4 = 0

OpenStudy (anonymous):

Now, this one has an x and y intercept

OpenStudy (anonymous):

I'll show you a site http://www.wolframalpha.com/input/?i=y%2B4%3D0 It shows where the equation meets with the x and y axis so you can tell your x and y coordinates

OpenStudy (anonymous):

But I have to show work

OpenStudy (anonymous):

@hedyeh99 I dont' understand :(

OpenStudy (anonymous):

It is confusing http://www.algebralab.org/lessons/lesson.aspx?file=algebra_lineareqintercepts.xml this might help